interest rate is 9% doubling time is 8 years if you invest $5000.00 what wil it grow to in 24 years?
a.$40,000
b.$15,000
c.$20,000
d.$30,000
do not understand how to fit this into a formula to get the answer please help have several more on this order to complete

Answer by stanbon(75887)

interest rate is 9% doubling time is 8 years
--------------------------------------------
A(t) = Ao*2^(t/8)
---------------------------
if you invest $5000.00 what will it grow to in 24 years?
A(24) = 5000*2^(24/8)
A(24) = 5000*2^3
A(24) = 5000*8
A(24) = $40,000
--------------------------------------------
Cheers,
Stan H.
===========================================
a.$40,000
b.$15,000
c.$20,000
d.$30,000